OpenStudy (anonymous):

Can someone help me with this? x -2 x^2+3x -10

OpenStudy (anonymous):

what do you want to do with it?

OpenStudy (anonymous):

Simplify the rational expressions. State any excluded values. Show your work.

OpenStudy (anonymous):

that's a polynomial expression. A rational expression is a fraction

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Other then that, that's all i have for information

OpenStudy (anonymous):

is that the exact way it's written? there's no numerator/denominator?

OpenStudy (anonymous):

2x/x^2+3x-10 sorry didn't include the division.

OpenStudy (anonymous):

like this? \[\frac{ 2x }{ x^2+3x-10 }\]

OpenStudy (anonymous):

Yes like that

OpenStudy (anonymous):

Usually you'd factor like this and something will cancel from the numerator and denominator. \[\frac{ 2x }{ (x+5)(x-2) }\] Since nothing canceled, the original is already simplified. For excluded values set the denominator equal to 0. \[(x + 5)(x - 2) = 0\] So you need to solve \(x+5=0\) and \(x-2=0\)

OpenStudy (anonymous):

Alright. x+5-5=0-5 x= -5 X-2+2=0+2 = 2

OpenStudy (anonymous):

yes, so -5 and 2 are the excluded values

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

So i'm finish @peachpi?

OpenStudy (anonymous):

yeah, unless they asked for something else

OpenStudy (anonymous):

Alright thanks :) bye.